# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

### v1

@@ 3.3. Query Transformation using GenAI
Fortunately, recent advancements in generative AI, particularly the emergence of the foundation models (e.g., DALL·E 3) based on guided-diffusion techniques (Zhang et al., 2023; Rombach et al., 2022; Ramesh et al., 2022), have enabled the generation of synthetic multi-modal tuples for complex prompts $p$ (natural language descriptions of the tuple to be generated). However, due to many reasons such as the randomized nature of the foundation models, the generated tuples [^1] Throughout this paper, we refer to these synthetic tuples as “guide tuples” that represent the natural language query in search space. We use these tuples to “guide” our system toward retrieving the most relevant results. may not be the optimal synthetic tuple $\mathbf{g}_{\varphi}$, but we can assume that the generated tuple is (semantically) in the vicinity of $\mathbf{g}_{\varphi}$.
Formally, we assume every tuple generated by the foundation model for the prompt $p=toPrompt(\varphi)$ lies in an $\varepsilon$ distance from $\mathbf{g}_{\varphi}$. Where $\mathbb{E}[\varepsilon]=0$. In other words for a sample tuple $\bar{\mathbf{g}}_{\varphi}$ we have,
$\mathbb{E}\left[\Delta(\mathbf{g}_{\varphi},\bar{\mathbf{g}}_{\varphi})\right]=0$ $\Rightarrow\mathbb{E}\left[\delta\left(\mathcal{E}^{o}(\mathbf{g}_{\varphi}),\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right)\right]=0$
$\Rightarrow\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$
@@
3.4. The Monte Carlo Algorithm
We are now ready to develop our algorithm: a Monte Carlo randomized algorithm based on repeated sampling.
Given a natural language query $\varphi$, and at least one foundation model $\mathcal{F}$, at the first step, we generate $m$ synthetic guide tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$. As mentioned before, we consider each guide tuple $\bar{\mathbf{g}}_{i}$ as an unbiased sample taken from a distribution with mean $\mathbf{g}_{\varphi}$.
Instead of relying on the distance estimations one vector representation, we view each embedding as a sample from the space possible transformation functions. Hence, we use multiple embedders to minimize the semantic distance estimation error. Let $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$ be the set of learned vector representations, considered for transformation. For each embedder $\mathcal{E}^{\ell}$, we assume $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$.
@@ 3.4. The Monte Carlo Algorithm
(3) $\bar{\delta}_{\varphi,i}=\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)$
Theorem 3.1.
@@
Proof.
Replacing $\bar{\delta}_{\varphi,i}$ with the right-hand side of Equation 3, we get
@@
Next, applying Chernoff bound, we get
@@
Therefore, applying the union bound, we get
@@
Using the estimated distances, we identify the $k$-nearest neighbors of $\varphi$ as
4. System Development
Figure 2. Needle Architecture
Having discussed the theoretical foundation of our system, we can now explain its architecture and different components in this section. While our GenAI-powered Monte Carlo method is modality agnostic, we will explain the development of our system for Image Retrieval. The overall architecture of Needle is illustrated in Figure 2. The architecture can be divided into two main parts: components involved in Preprocessing [§ 4.1, § 4.2, § 4.3] and Inference [§ 4.4, § 4.5, § 4.6, § 4.7].
4.1. Preprocessing
The preprocessing phase involves a one-time operation to create vector representations, also known as embeddings, of the input dataset using a set of predefined embedders. Since images are typically regarded as being strongly associated with their representing objects(Zhao et al., 2019), embedding models may underperform when applied to images containing numerous small objects (Liu et al., 2021b).
To address this limitation, we employ a tiling strategy that divides the input image into multiple tiles based on the density of objects in each region (§ 4.2).
Each of these tiles is then processed by the embedders, and the resulting embeddings are stored on disk. This process can be executed in parallel for each embedder, with performance significantly enhanced by using multiple GPUs to handle tile embedding.
4.2. Smart Tiling
Figure 3. An Illustration of Smart Tiling on an image with lots of objects
Algorithm 1 Smart Tiling
1:image, d, objects_borders, min_size
2:image tiles
3:$(h,w)\leftarrow\text{size of }image$
4:tiles [$\leftarrow$[]← [ ]
5:function CountObjects(x, y, w, h)
6:    return number of unique objects in the region $(x,y,w,h)$
7:end function
8:
9:function Split(x, y, w, h)
10:    num_objects $\leftarrow$ CountObjects(x, y, w, h)
[…]
12:         Append((x, y, w, h)) to tiles
13:    else
14:         Compute half_w, half_h
15:         Calculate objs_left, objs_right, objs_top, objs_bottom
16:         if vertical split is preferred & aspect ratio is reasonable then
17:             Split(x, y, half_w, h) $\triangleright$ Split vertically
18:             Split(x + half_w, y, half_w, h)
19:         else
20:             Split(x, y, w, half_h) $\triangleright$ Split horizontally
21:             Split(x, y + half_h, w, half_h)
22:         end if
23:    end if
24:end function
25:
26:Split(0, 0, w, h)
27:return tiles
To capture all the details within an image, it is important to focus on densely populated regions. Therefore, we should split each image into different tiles based on the density. In addition, most embedders are trained on square images (He et al., 2016; Simonyan and Zisserman, 2014; Szegedy et al., 2017) and perform optimally when applied to such shapes (Shahnaz and Mollah, 2022; Azulay and Weiss, 2019), which means excessively tall or narrow tiles should be avoided. In order to achieve these two goals, inspired by (Finkel and Bentley, 1974), we propose our Smart Tiling procedure. We employ an edge detection technique (Canny Edge Detection (Canny, 1986)) to identify object boundaries within the image and recursively split the image to balance the distribution of objects across tiles as effectively as possible. The objective of this algorithm is to divide the image into smaller tiles, where each tile contains no more than a specified number of objects $d$, while maintaining a balanced aspect ratio.
The algorithm recursively splits the image by first checking if a region contains more objects than the allowed threshold $d$. If this condition is met, the region is split either vertically or horizontally. The algorithm calculates the object distribution for potential splits and selects the one that provides the most balanced distribution of objects across the two resulting tiles. To avoid generating disproportionate tiles, the algorithm also checks the aspect ratio and favors splits that maintain near-square dimensions. This recursive process continues until each tile satisfies the object limit constraint or it reaches the minimum acceptable width/height, at which point the tiles are finalized. Algorithm 1 presents the pseudocode for our Smart Tiling approach.
4.3. Vector Store
Once the raw image data is processed into vectors, it is indexed within the vector store. The vector store provides a maximum inner product search interface with low latency, which is crucial since users expect prompt results from the system. In our implementation, we employ hnswlib, which offers approximate maximum inner product search capabilities. This is consistent with the functionality provided by most vector stores.
4.4. Query Refinement
Having discussed the components involved in Preprocessing, we now explore the components used during Inference. After receiving the query in natural language text, Needle refines the input query into a suitable prompt. The first step is to remove punctuation and redundant characters from the query. Additionally, Needle supports more advanced query refinement techniques.
[…]
In addition, given that current Foundation Models have limitations in fully understanding some types of prompts (Liu et al., 2022; Feng et al., 2022; Lee et al., 2023; Liu et al., 2024), our system can employ various query refinement techniques (Liu et al., 2021a; Liu and Chilton, 2022; Wang et al., 2023; Witteveen and Andrews, 2022) to paraphrase and reformat the input query. These transformations help better align the generated query tuples with the original intent. These techniques help overcome prompt misinterpretations and improve the relevance of generated results.
4.5. Foundation Model and Guide Image Generation
After refining the query, We need to generate guide tuples representing the query using foundation models. Needle can use multiple (black-box) foundation models to generate guide images, with the number of guide images fine-tuned as a hyperparameter, denoted as $m$. Foundation models are capable of producing images with varying levels of quality and resolution. However, for the purposes of Needle, high-quality realistic images are unnecessary to achieve accurate search results. Lower-resolution images are sufficient, provided that they closely resemble the input query. In our experiments (§ 6.3), we will explore the impact of $m$, the size of the generated images, and the overall quality of the foundation model in the final results.
4.6. User Interface and Feedback Process
Figure 4. Needle User Interface
The current implementation of Needle offers a web-based user interface that allows users to search based on their queries (Figure 4). Users can interact with the system in two modes: Feedback mode (on) or No-Feedback mode (off). When Feedback mode is enabled, the guide tuples are presented to the user prior to the search process. Users can evaluate the quality and relevance of the generated guide tuples with respect to their query, which can assist Needle in retrieving more relevant images. In No-Feedback mode, Needle automatically identifies and discards outlier tuples from the generated images without user intervention (§ 5.2).
4.7. Embedders and Rankings
After generating the guide images, Needle utilizes predefined embedders to convert the guide tuples into vector representations and performs a k-NN lookup on the vector stores (§ 4.3). The score of each image is determined by the maximum score among all of its tiles (with the entire image also treated as a tile). Subsequently, the results are aggregated using our aggregation mechanism (§ 5.1) to produce a final ranking, which is then presented to the user.
5. Practical Optimization
In this section, we discuss our practical optimizations to improve the efficiency of our system.
First, we address the issue of aggregating results from multiple embedders into a unified ranking (§ 5.1), followed by a discussion on detecting anomalies in generated synthetic guide tuples (§ 5.2).
5.1. Rankings Aggregation and Embedder Trust Mechanism

### v2

@@ 3.1 Query Transformation using GenAI
Fortunately, recent advancements in generative AI, particularly the emergence of the foundation models (e.g., DALLE 3) based on guided-diffusion techniques [80, 57, 54], have enabled the generation of synthetic multi-modal tuples for complex prompts $p$ (natural language descriptions of the tuple to be generated). However, due to many reasons such as the randomized nature of the foundation models, the generated tuples [^4] Throughout this paper, we refer to these synthetic tuples as “guide tuples” that represent the natural language query in search space. We use these tuples to “guide” our system toward retrieving the most relevant results. may not be the optimal synthetic tuple $\mathbf{g}_{\varphi}$, but we can assume that the generated tuple is (semantically) in the vicinity of $\mathbf{g}_{\varphi}$.
Formally, we assume every tuple generated by the foundation model for the prompt $p=toPrompt(\varphi)$ lies randomly within a ($d_{\xi^{o}}$-dimensional) ball centered at $\mathbf{g}_{\varphi}$ with the radius $\varepsilon$. Therefore, for a sample tuple $\bar{\mathbf{g}}_{\varphi}$ we have,
$\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$
@@
3.2 The Randomized Algorithm
We are now ready to develop our Monte Carlo randomized algorithm based on repeated sampling.
Given a natural language query $\varphi$, and at least one foundation model $\mathcal{F}$, at the first step, we generate $m$ synthetic guide tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$. As mentioned before, we consider each guide tuple $\bar{\mathbf{g}}_{i}$ as an iid (independent and identically distributed) sample from a distribution with mean $\mathbf{g}_{\varphi}$. Furthermore, as mentioned earlier, instead of relying on the distance estimations based on one vector representation, we view each embedder $\mathcal{E}^{\ell}$ as a sample from the space possible embedders. Hence, we use a set $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$ of embedders to minimize the semantic distance estimation error. For each embedder $\mathcal{E}^{\ell}$, we assume $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$.
@@ 3.2 The Randomized Algorithm
$\hskip 56.9055pt\bar{\delta}_{\varphi,i}=\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)$ (4)
We propose following theorem to show that with a high probability the distance estimations by Equation 4 are near-optimal.
Theorem 1
 [^5] The proof is provided in Appendix B.
@@
$\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}.$
Following Theorem 1, we can use the estimated distances to identify the $k$-nearest neighbors of $\varphi$ as
3.3 Practical Optimizations
In addition to the theoretical results, we implement several practical optimizations to enhance our system’s performance and efficiency. Due to space constraints, we provide a concise summary here, with full details elaborated in Appendix C. First, we consider aggregating the rankings based on embedders’ trust scores, which are learned on the fly. Since embedders are trained on different datasets and employ different architectures, they may perform differently in practice. Therefore, we assign a performance (trust) weight to each embedder conditioned on each topic. We propose an algorithm that dynamically adjusts the embedders’ weights based on their performance across various topics using users’ feedback, thereby optimizing their contribution to the final aggregated ranking (Appendix C.1). To improve the robustness, we incorporate an outlier detection mechanism to discard potentially misaligned or corrupted generated images from the guide set (Appendix C.2). Furthermore, we enhance inference speed through two parallel mechanisms: a caching system that leverages the previously issued queries and acquired metadata to expedite similar query responses (Appendix C.3), and a query classifier that distinguishes simple from hard queries, allowing direct retrieval for simple cases without requiring guide image generation (Appendix C.4).
4 System Overview
In the preceding section, we presented the theoretical foundations of our approach. In this section, we detail how these principles form the design of our image content retrieval system. Although our GenAI-powered Monte Carlo method is modality-agnostic, our focus here is on its application to image retrieval. Figure 2 illustrates the overall architecture of Needle, which is organized into two primary components: Preprocessing and Inference. [^6] System development details is provided in Appendix D
Preprocessing: The preprocessing phase involves a one-time operation to create vector representations of the input dataset using a set of predefined embedders $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$. Once the raw image data is processed into vectors, it is indexed within the vector store. The vector store provides a maximum inner product search interface with low latency, which is crucial since users expect prompt results from the system.
Inference: At the query time, we need to generate guide tuples representing the query using foundation models. Needle can use multiple (black-box) foundation models to generate guide images, with the number of guide images fine-tuned as a hyperparameter ($m$). Different foundation models might produce images with varying levels of quality and realism. However, for the purposes of Needle, high-quality, realistic images are unnecessary to achieve accurate search results. As a result, lower-quality images are sufficient, as long as they closely resemble the input query. In our extended experiments in Appendix E.4, we will explore the impact of different configurations in the image generation process, like the size of the generated images, the number of guide images, the count, and the overall quality of the foundation models in the final results.
After generating the guide images, Needle uses predefined embedders to transform guide tuples into vector representations and conducts a k-NN lookup for each generated image across all embedders. The results are then aggregated into a final ranking, where each embedder’s contribution is calculated according to its current weight. This weight reflects the system’s trust in the performance of the corresponding embedder. While dynamic adjustment of these weights based on user feedback is supported (Appendix C.1), the default configuration assumes such feedback is unavailable and uses fixed predefined weights, which can be derived from external performance indicators like timm leaderboard scores [^7] https://huggingface.co/spaces/timm/leaderboard. This weighting mechanism is general and fully configurable by the user.
Figure 2: Needle Architecture
5 Experimental Evaluation
In this section, we evaluate several key aspects of Needle to demonstrate its capabilities and to evaluate its performance. We use multiple benchmark datasets and several baselines for this purpose. In the following, we first detail our experiments setup (§ 5.1), followed by a proof-of-concept analysis (§ 5.2) that demonstrates Needle’s efficacy in text-to-image retrieval across diverse object detection (§ 5.2.1) and complex natural language query benchmarks (§ 5.2.2). Subsequently, we conduct an ablation study in Appendix E.4 to assess the impact of hyperparameter variations on Needle’s performance. Finally, in Appendix E.5 we provide an end-to-end case study, involving human participants’ qualitative feedback on Needle’s responsiveness to arbitrary user-generated queries.
5.1 Experiments Setup
Baselines: We evaluate our method using several prominent vision-language models as baselines, representing different architectural and training paradigms. These include well-established contrastive learning models such as CLIP [53] and ALIGN [25], the unified multimodal approach of FLAVA [60], and the effective decoupled pipeline combining BLIP for captioning with MiniLM for text embedding [35, 72]. Each of these models offers a unique perspective on learning joint image-text representations, and together they provide a comprehensive set of comparisons for our work. Detailed descriptions of these models and their specific configurations used in our experiments are deferred to Appendix E.1.
Datasets: We evaluate Needle using a comprehensive set of datasets, categorized into those focusing on object detection and those involving complex natural language queries. The object detection benchmarks include Caltech256 [21], MS COCO [38], LVIS [22], and BDD100k [79]. For complex natural language queries, we utilize COLA [56], Winoground [63], NoCaps [2], and SentiCap [44]. These datasets collectively provide diverse challenges, from identifying specific objects to understanding nuanced language and handling novel concepts or sentiments, essential for benchmarking zero-shot retrieval performance. Detailed descriptions of the datasets are provided in Appendix E.2.
Evaluation Metrics: We employ distinct evaluation metrics tailored to the characteristics of our two benchmark types: object detection and complex natural language queries. For object detection benchmarks, we measure performance using Mean Recall at $k$ (R@k), Precision at $k$ (P@k), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR). For complex natural language query experiments, we report both MRR and Pairing Accuracy (PAcc). Detailed definitions and formulas for these evaluation metrics are provided in Appendix E.3.
Embedders: We use all or a subset of embedders listed in Appendix (Table 3) in our experiments.
Hardware Configuration: All experiments were conducted on two servers, each equipped with 32 GB of memory, a 12-core CPU, and two Tesla T4 GPUs, running Ubuntu 22.04 LTS.
Foundation Models and the Monetary Cost: For image generation in our experiments, we employed several state-of-the-art foundation models, namely DALL-E [^8] https://openai.com/index/dall-e-3/, ImagenV3-fast [^9] https://replicate.com/google/imagen-3-fast, Flux-Schnell [^10] https://replicate.com/black-forest-labs/flux-schnell, and RealvisXL-v3.0-turbo [^11] https://replicate.com/adirik/realvisxl-v3.0-turbo. These models were chosen for their complementary strengths in generating high-quality images. The total monetary cost incurred for utilizing these models was \$257.53 (USD).
5.2 Proof of Concept
Figure 3: Illustration of Complex Natural Language Queries extracted from NoCaps [2], Guide images, and CLIP vs. Needle results.
5.2.1 Task 1: Object Detection
Table 1: Zero-shot retrieval performance on object detection datasets. For each dataset, we report R@10, P@10, MAP, and MRR. Each cell shows two numbers (All / Hard).
Caltech256 COCO LVIS BDD
Baseline R@10 P@10 MAP MRR R@10 P@10 MAP MRR R@10 P@10 MAP MRR R@10 P@10 MAP MRR
CLIP 0.926 0.150 0.926 0.150 0.939 0.181 0.952 0.193 0.934 0.400 0.934 0.400 0.952 0.477 0.988 1.000 0.177 0.093 0.167 0.083 0.168 0.078 0.316 0.224 0.660 0.033 0.660 0.033 0.670 0.005 0.714 0.048
ALIGN 0.941 0.375 0.941 0.375 0.947 0.398 0.961 0.541 0.944 0.800 0.944 0.800 0.960 0.895 0.981 1.000 0.215 0.145 0.201 0.130 0.207 0.129 0.379 0.306 0.560 0.000 0.560 0.000 0.573 0.003 0.704 0.014
FLAVA 0.882 0.258 0.882 0.258 0.903 0.306 0.949 0.491 0.924 0.100 0.924 0.100 0.941 0.281 0.963 1.000 0.185 0.109 0.172 0.097 0.180 0.099 0.321 0.241 0.660 0.033 0.660 0.033 0.698 0.036 0.725 0.083
BLIP + MiniLM 0.826 0.317 0.826 0.317 0.838 0.372 0.880 0.408 0.941 0.700 0.941 0.700 0.951 0.698 0.975 1.000 0.180 0.115 0.177 0.111 0.179 0.107 0.332 0.260 0.600 0.167 0.600 0.167 0.610 0.144 0.725 0.333
[…]
We begin our experiments by evaluating the performance of Needle against the baselines on object detection benchmarks, utilizing the Caltech256, COCO, LVIS, and BDD datasets. For each object in these benchmarks, we formulate queries to retrieve them. For Needle, we employed three image generation engines (Flux-Schnell, RealVisV3, and ImagenV3-fast) to generate three images per engine, resulting in a total of nine guide images per query, all at MEDIUM quality. Additionally, we utilized the ensemble of embedders detailed in Table 3. Following the methodology outlined in Appendix E.1, we considered queries with a CLIP Average Precision below 0.5 as the “hard set”.
Table 1 presents the performance of Needle against the baselines across R@10, P@10, MAP, and MRR metrics. Notably, Needle demonstrates superior performance compared to all baselines on both the easy and hard sets. Specifically, for the hard set, Needle achieves MAP improvements of 73%, 10%, 93%, and 10% over the second-best baseline on Caltech256, COCO, LVIS, and BDD datasets, respectively. These results underscore the substantial advantage of using synthetic data for image retrieval over traditional contrastive learning and image-to-text approaches. Furthermore, Figure 3 provides a visual illustration of the performance of Needle (v.s. CLIP) and the synthetic guide images generated for each query.
5.2.2 Task 2: Complex Natural Language Queries
Table 2: Zero-shot retrieval performance on complex natural language queries benchmarks.
Cola Winoground SentiCap NoCaps
Baseline PAcc MRR PAcc MRR MRR MRR
CLIP 0.578 0.246 0.519 0.426 0.464 0.573
ALIGN 0.591 0.301 0.554 0.501 0.555 0.704
FLAVA 0.615 0.336 0.574 0.482 0.546 0.658
BLIP + MiniLM 0.449 0.195 0.485 0.330 0.331 0.398
[…]
Experiments involving complex natural language queries evaluate the system’s ability to handle challenging language and visual nuances across different retrieval tasks. We define two types of evaluations:
First, the Pairing Task assesses the system’s capability to correctly match a caption to its corresponding image when presented with two highly similar images and two corresponding captions that differ only in subtle details. Given four possible image-caption pairings, random chance achieves 0.25 Pairing Accuracy (PAcc). This task is evaluated using the Winoground [63] and COLA [56] datasets (see Figure 6 for examples from COLA illustrating the challenges of this setup).
Second, the Full-set Retrieval Task measures the system’s performance when retrieving a specific image from the entire dataset using a single, potentially complex, query. For this task, we report the Mean Reciprocal Rank (MRR). This task is evaluated on the Winoground [63], COLA [56], SentiCap [44], and NoCaps [2] benchmarks, as they feature challenging captions that require nuanced understanding for successful retrieval from a large pool.
Table 2 presents the Pairing Accuracy and Mean Reciprocal Rank (MRR) metrics for different aforementioned baselines. As illustrated in the table, Needle outperforms the baselines in both PAcc and MRR. It is important to note that current foundation models still exhibit limitations in generating images that are fully aligned with an input query. In particular, our investigation reveals that these models frequently fail to produce images with the correct compositional ordering. We anticipate that further advancements in foundation models will further enlarge the performance gap between Needle and the baselines. It is also noteworthy that the BLIP+MiniLM baseline performed significantly worse than the other baselines. Initially, we anticipated that this baseline would serve as a strong competitor in handling complex natural language queries, given its design to capture compositional and relational attributes between objects in images; however, the empirical results did not corroborate these expectations. In contrast, FLAVA demonstrated exceptionally strong performance in these tests.
6 Conclusion
Following this method, We developed Needle, an open-source database dedicated to delivering robust, efficient, and high-performance image retrieval [^12] Due to the space constraints, the limitations of our system are discussed in Appendix A.. Our comprehensive experiments demonstrated that our system significantly outperformed the state-of-the-art techniques in image retrieval.
While our proposed technique is not limited to image datasets, its development for other modalities, such as audio and video, requires foundation models that are currently not publicly available. Upon the availability of such models, we aim to extend the scope of Needle to a wider range of modalities in our future work. We appreciate and highly value future contributions from the open-source community—whether through feedback, bug reports, feature requests, or direct code contributions—to help us further enhance and refine the system.
APPENDIX
Appendix A Discussions and Limitations
Reliance on Existing Models: Needle relies on Foundation Models for guide tuple generation and on embedders for generating vector representations of images, which may have their own capabilities, limitations, and inherent biases. This reliance presents both strengths and weaknesses in Needle’s architecture. The strength lies in its ability to improve performance by upgrading the internal Foundation Models and embedders. On the other hand, though, the limitations of these models also limit Needle. To mitigate this issue, Needle supports multiple Foundation Models and embedders, enabling it to draw from a broader range of knowledge and reduce the impact of individual model limitations.
Preprocessing Efficiency: During the preprocessing phase, we employ a predefined set of embedders to generate vector representations of images. Needle processes the images in parallel while utilizing the available GPU resources to distribute the workload efficiently. Moreover, Needle supports multiple operational modes upon installation. For instance, the fast mode utilizes only two embedders which offers preprocessing speeds that are comparable to those of state-of-the-art methods.
Inference Efficiency: Needle’s inference consists of three main sequential operations: (a) generating guide tuples, (b) k-NN lookup, and (c) ranking aggregation. Theoretically, for retrieving $k$ results from a dataset of size $n$ with constants $m$ (number of guide images) and $\ell$ (number of embedders), Step (a) is a constant-time operation. Step (b)’s time complexity depends on the underlying vector store algorithm, and in the current implementation, we leverage HNSW data structure index [42], which provides k-NN lookup time complexity of $O(k\log(n))$, while Step (c) uses Fagin’s instance-optimal TA algorithm for rank aggregation [17].
In practice, however, the bottleneck lies in Step (a), generating guide tuples. For instance, in all of our experiments with up to $n=120K$ images, steps (b) and (c) took less than a second, dominated by the image generation time. Fortunately, the image generation time is independent of the dataset size $n$ or the output size $k$, and only relies on the number of guide images $m$. Moreover, the generation process can run in parallel, significantly improving performance. Needle’s current implementation supports both on-device and online image generation models. For optimal performance, we recommend using fast on-device generation models to eliminate network latency. Moreover, as demonstrated in our ablation study, Needle maintains high accuracy even with lower-quality, lower-resolution images. Therefore, employing fast, lower-quality foundation models can significantly accelerate the overall process. In addition to these, the builtin Query Complexity Classifier (C.4) enables Needle to generate images only when the query deemed complex for preliminary methods, which reduces the average inference time drastically compared to using Needle for all queries.
Extensions to Other Modalities: The core idea behind Needle is adaptable to other modalities beyond images, including audio and video. However, due to the current limitations in publicly available foundation models, our focus in this paper was on the image data. As foundation models for these additional modalities continue to improve and become more accessible, we plan to extend Needle to support them. This will allow Needle to operate across multiple data types, unlocking further potential for handling complex, multi-modal queries in the near future.
Appendix B Proof of Theorem 1
Theorem 1 Let $\delta_{\varphi,i}=\delta\left(\vec{q}(\varphi),\vec{v}^{o}_{i}\right)$, $\forall t_{i}\in\mathcal{D}$. Given a positive value $\gamma$,
@@ Appendix B Proof of Theorem 1
Proof: Replacing $\bar{\delta}_{\varphi,i}$ with the right-hand side of Equation 4, we get
@@
Next, applying Chernoff bound [49], we get
@@
Therefore, applying the Union bound [49], we get
@@
Appendix C Practical Optimization
C.1 Rankings Aggregation and Embedder Trust Mechanism